OpenStudy (saifoo.khan):

What are? Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem.

OpenStudy (saifoo.khan):

1014 = x^3 - 7x^2

OpenStudy (saifoo.khan):

1014 = x^3 - 7x^2 I have to solve this equation by these three methods.

OpenStudy (amistre64):

rational roots is a way to narrow down the pool of trial and error options. ax^2 +bx + c = 0 x^2 +bx/a + c/a = 0 the last term is important in that if this is going to come out pretty; it need to factor as c/a

OpenStudy (amistre64):

so we take the: factors of "c" -----------; plus and minus as our options factors of "a"

OpenStudy (amistre64):

Descartes' Rule of Signs is a way to figure out how many + or - zeros are possible

OpenStudy (saifoo.khan):

in this case, the factors of 1014 ---------------- factors of 1

OpenStudy (saifoo.khan):

Sorry i have no idea how to do them.

OpenStudy (amistre64):

the factors of 1 are 1 so that bottom is a bit moot

OpenStudy (amistre64):

1014 factors into ....

OpenStudy (amistre64):

1,1014 2,507 .. etc

OpenStudy (saifoo.khan):

then?

OpenStudy (saifoo.khan):

1 | 2 | 3 | 6 | 13 | 26 | 39 | 78 | 169 | 338 | 507 | 1014 -------------------------------------------------------------- 1

OpenStudy (amistre64):

then you have 2 options; use your pool as trial and errors in your f(x) and whatever pops out a zero is good

OpenStudy (amistre64):

well, the other option is really the same, just a different format

OpenStudy (saifoo.khan):

but what's the method? i have to show! :(

OpenStudy (amistre64):

what the f(x) ?

OpenStudy (saifoo.khan):

f(x) = x^3 - 7x^2 -1014

OpenStudy (amistre64):

well, it a cubic; which tells us there are 3 zeros, some might be multiples tho

OpenStudy (saifoo.khan):

i know the answers, they are: 13, 13 and 6.

OpenStudy (saifoo.khan):

they are the sides of a cuboid.. i had to make an equation using these... So, i used the formula of volume..

OpenStudy (amistre64):

f(1) = 1 - 7 -1014 = not a zero f(2) = 8 - 28 -1014 = not a zero etcc

OpenStudy (saifoo.khan):

they will be 13, 13 and 6.

OpenStudy (saifoo.khan):

how to solve it by factor theorm?

OpenStudy (amistre64):

f(6) = 6^3 - 7.6^2 -1014 f(6) = 216 - 7.36 -1014 f(6) = 216 - 252 -1014 f(6) = 216 - 252 -1014; you sure its a 6?

OpenStudy (amistre64):

13 is good

OpenStudy (saifoo.khan):

Let me explain, i had to make an equation in terms of x.. with the volume of a cuboid. And i have to use x+something & x + something for width and heights. So, Volume = L * B * H V = 13 * 13 * 6 V = 1014

OpenStudy (amistre64):

factor thrm i got no idea what the definition would be, but id assume its factoring it

OpenStudy (saifoo.khan):

Now, Length = x breath = x height = (x - 7)

OpenStudy (saifoo.khan):

1014 = l * b* h 1014 = x * x * (x - 7)

OpenStudy (saifoo.khan):

that's how i made the equation.

OpenStudy (amistre64):

ok, and x^3 - 7x^2 = 1014 x^3 - 7x^2-1014 = 0 right?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (saifoo.khan):

So now?

OpenStudy (amistre64):

so now, factor thrm is just factoring it back out to: x^2(x-7) i believe

OpenStudy (saifoo.khan):

but i dont get 6 then! :(

OpenStudy (amistre64):

or rather: (x-13)(x^2+6x+78) = 0

OpenStudy (saifoo.khan):

Great^

OpenStudy (saifoo.khan):

what about another 13 and 6?

OpenStudy (amistre64):

x = 13, not 6

OpenStudy (saifoo.khan):

how will will i get the height and the width?

OpenStudy (amistre64):

you found x; use that to find the height and the width

OpenStudy (amistre64):

width = x length = x height = x-6

OpenStudy (amistre64):

since x=13 .... plug it in

OpenStudy (saifoo.khan):

owow^

OpenStudy (saifoo.khan):

So that's factor therom.. what about others? :/

OpenStudy (amistre64):

the others is the finding a pool of options to trial and error with

OpenStudy (saifoo.khan):

Descartes' Rule of Signs <------------ i never heard of that. :(

OpenStudy (amistre64):

rational roots pops out f(13) = 0 , and 13 is one of the factors of 1014

OpenStudy (amistre64):

Ds rule of sign is simply a way to measure the max posibble number of positive or negative zeros

OpenStudy (amistre64):

f(x) = x^3 -7x^2 -1014 ----> ----------> 1 0 there is only 1 sign change which suggests a maximum of 1 positive zero f(-x) = -x^3 -7x^2 -1014 -----------------> 0 there are NO sign changes, so there are a maximum of 0 negative roots

OpenStudy (anonymous):

by factor theorem f(13)=0 so (x-13) is factor of x^3-7x^2-1014 now divde x^3-7x^2-1014 by x-13 qutient is x^2-6x-78 so equation becomes (x-13)(x^2-6x-78)

OpenStudy (saifoo.khan):

Oh! so that's ds... that was short. why u made arrows below it? But the sign of x^3 changed! :O

OpenStudy (amistre64):

to show the progression from sign change to sign change and count them

OpenStudy (amistre64):

f(-x) imples a negative value for a zero

OpenStudy (amistre64):

f(-x) would pop out: -x^2 -7x -1014 which produces NO changes in sign thru the equation

OpenStudy (saifoo.khan):

i dont get it. :(

OpenStudy (saifoo.khan):

sorry.

OpenStudy (amistre64):

spose g(x) = x^2 +6x +4; this would have 2 roots since its a parabola that dips under the x axis right? g(x) = x^2+6x+4 ----------> no change in sign; max number of +roots, none g(-x) = x^2 -6x +4 --->--->---> 1 1 0 there are 2 sign changes when x is negative; telling us there is a max possibility of 2 -roots

OpenStudy (amistre64):

it just so happens it has 2 -roots: http://www.wolframalpha.com/input/?i=x^2%2B6x%2B4

OpenStudy (amistre64):

what the root values are? Ds sign rule doesnt say

OpenStudy (saifoo.khan):

ok, so for my problem, i should copy down this thing? f(x) = x^3 -7x^2 -1014 ----> ----------> 1 0 there is only 1 sign change which suggests a maximum of 1 positive zero f(-x) = -x^3 -7x^2 -1014 -----------------> 0 there are NO sign changes, so there are a maximum of 0 negative roots

OpenStudy (amistre64):

If your being graded based upon material specific methods; no. If you are being graded upon how well you understand the concepts, then yes

OpenStudy (saifoo.khan):

methods. :(

OpenStudy (amistre64):

then youll have to render the appropriate material specific accounting for it. Which should be similar

OpenStudy (saifoo.khan):

or what do u adivce? if i make another equation and take different sizes, would that be easier?

OpenStudy (amistre64):

it isnt the equations that are bad; its simply the method by which you are to be graded by that is the issue here. And regardless of what your equation is, the method proscribed by your material is what you are graded by.

OpenStudy (amistre64):

without having read and studied what the author of your material does, I have no way of determining a suitable "gradable" response.

OpenStudy (saifoo.khan):

Alright, thanks.. if i stuck anywhere i will ask you! :)